OpenStudy (the_puppet1):

Simplify: -(n + 2) +(n - 2) A. 4n - 8 B. 4n + 8 C. 2n - 8 D. 2n + 8

OpenStudy (the_puppet1):

@reemii

OpenStudy (the_puppet1):

Plz help me

OpenStudy (the_puppet1):

@carolina4567

OpenStudy (carolina4567):

Any Ideas ?

OpenStudy (the_puppet1):

I think it's C

OpenStudy (jpharper12345678cars):

wait you should use mathway.com they have a thing that will give you the answer to this kind of stuff

OpenStudy (jpharper12345678cars):

you have to make sure its under simplyfy

OpenStudy (reemii):

if the expression is \(-(n+2) + (n-2)\), the answer is not in the options. Because \(-(n+2)+(n-2)=-n-2+n-2=-4\). Can you check the expression to simplify?